OpenStudy (anonymous):

“Ralphie has 23 more toys than his little brother Joey. If they were to get all of their toys out and lay them on the floor, there would be 384 toys on the floor.”

OpenStudy (anonymous):

this is the equation r-j=23 r+j=384

OpenStudy (whpalmer4):

Are you supposed to find the number of toys each kid owns?

OpenStudy (anonymous):

here are the answer choices , sorry A. The system of equations the student used is correct. B. The top equation in the system is incorrect. C. The bottom equation in the system is incorrect. D. Both equations in the system are incorrect.

OpenStudy (whpalmer4):

Well, what do you think?

OpenStudy (anonymous):

b ?

OpenStudy (whpalmer4):

Why do you think that?

OpenStudy (anonymous):

i dont know . . i guess because it says ralphie has more 23 more toys than his little brother Joey . .

OpenStudy (whpalmer4):

if ralphie has r toys, how many does joey have? isn't that just j = r-23?

OpenStudy (anonymous):

r - j = 23, r + j = 384 is correct. The numbers are wrong. \[\left\{r\to \frac{407}{2},j\to \frac{361}{2}\right\} \]The answers should be integers.

OpenStudy (anonymous):

thank you (:

OpenStudy (whpalmer4):

or r = j + 23, or r - j = 23?

OpenStudy (whpalmer4):

I hope @robtobey didn't really need Mathematica to solve that system and spot the problem :-)

OpenStudy (anonymous):

{r - j = 22, j + r = 384} yields {r = 203, j = 181}

OpenStudy (anonymous):

so the answe is A ?

OpenStudy (anonymous):

@whpalmer4 I never sign onto this site without it. Mathematica never makes mistakes.

OpenStudy (whpalmer4):

If you just add the 2 equations given together: r-j = 23 r+j = 384 --------- 2r = 407 Who owns half a toy?

OpenStudy (anonymous):

ok . .

OpenStudy (whpalmer4):

Doesn't tell you which equation is incorrect, but does alert you that there is a problem, and would be a natural first step in solving this particular set of equations.

OpenStudy (anonymous):

your right i get it now , thank you